OpenStudy (anonymous):

Find the value of the square root of 8 over 3 times the square root of 4. in simplest form. 6 2 over 3. the square root of 2 over 3. 3 times the square root of 2.

OpenStudy (anonymous):

@franciscanmonk can you please help?

OpenStudy (anonymous):

Somebody lease help

OpenStudy (anonymous):

okay let me see if I understand .. the top is sqrt8 which can reduce .. sqrt(2*2*2) =2sqrt2 ... the bottom is 3 * sqrt4 =3*2=6 .. you can reduce that all sqrt(2)/3

OpenStudy (anonymous):

Wouldnt it just be 2 over 3 then?

OpenStudy (anonymous):

sqrt(8)/3sqrt(4) 2sqrt(2)/3(2) sqrt(2)/3 sqrt(2)/3 <-- answer. Im kinda doing an assignment too.

OpenStudy (anonymous):

\[\frac{\sqrt{8}}{3\times \sqrt{4}}=\frac{2\times \sqrt{2}}{2\times3}=\frac{\sqrt{2}}{3}\]

OpenStudy (anonymous):

tk

OpenStudy (anonymous):

@franciscanmonk thanks man

OpenStudy (anonymous):

No probs.